OpenStudy (anonymous):

The first term of an arithmetic sequence is -3 and the fifteenth term is 53. What is the common difference of the sequence?

OpenStudy (aum):

The \(n^{th}\) term of an arithmetic sequence is: \(a_n = a_1 + (n-1)d\) where \(a_1\) is the first term and \(d\) is the common difference. Here, \(a_1 = -3;~~n = 15;~~a_{15} = 53\). Plug in and find d.

OpenStudy (anonymous):

where does a15 go?

OpenStudy (aum):

The left hand side. when n = 15, \(\large a_n \text{ becomes } a_{15}\)

OpenStudy (aum):

\[\large a_n = a_1 + (n-1)d \\ \large a_1 = -3;~~n = 15;~~a_{15} = 53 \\ \large 53 = -3 + (15 - 1)d \\ \large 53 + 3 = 14d \\ \large 56 = 14d \\ \large d = 56/14 = 4 \]

OpenStudy (anonymous):

Thanks!!

OpenStudy (aum):

You are welcome.